# Adaptation of Agentic AI

### v2

@@ 4.1.2 RLVR-Based Methods
Formal theorem proving [68, 69, 70, 71, 72, 73, 74, 75, 76] has emerged as a canonical domain for RLVR under the A1 paradigm, as proof assistants provide ground-truth, tool-execution–signaled feedback at every step. In this setting, the agent proposes one or more tactics (i.e., proof steps), a formal proof checker (the tool) deterministically verifies their validity, and the resulting validated proof-state transition is returned to the agent. This verification outcome—e.g., whether a tactic is accepted, whether it advances the proof state, or whether a complete proof is achieved—serves directly as a verifiable reward signal for policy optimization. Compared to code-execution RLVR, where unit tests may be sparse or incomplete, theorem proving offers step-wise semantic verification with minimal ambiguity, enabling denser rewards and substantially easing long-horizon credit assignment. Recent systems such as AlphaProof [77] (Nature 2025), DeepSeek-Prover-V2 [78] (ICLR 2025), Kimina-Prover [72], and Leanabell-Prover-V2 [79] leverage this verifier feedback to train multi-step proof search policies via reinforcement learning, while a complementary line of work augments the native proof checker feedback with auxiliary guidance signals to prioritize trajectories, shape exploration, or stabilize optimization on top of verifier-grounded rewards [80, 81, 82, 83, 84]. While RLVR is well suited for learning proof strategies under a fixed prover snapshot, formal theorem proving also highlights a broader adaptation challenge: formal libraries (e.g., mathlib [85]) and large, actively evolving formalization projects [86] built by the Lean community grow continuously, expanding the available premise space. Addressing this non-stationarity often requires complementary continual or low-resource adaptation mechanisms beyond pure RLVR, which we discuss in Continual Adaptation8.2).
Multi-tool reasoning systems incorporated multiple tools in sequential or compositional pipelines, which used environment feedback to guide action selection. Router-R1 [87] (NeurIPS 2025) trained a policy language model to coordinate multiple large language models through a RL framework that formulates multi-round routing and aggregation as a sequential decision process. During training, the policy LLM learns to alternate between internal reasoning and external model selection, dynamically invoking different LLMs from a routing pool to solve complex tasks. FTRL [88] proposed an automated strategy for constructing tool-use training environments through a multi-stage pipeline, enabling the creation of diverse and comprehensive training settings without external toolsets. Building on these environments, FTRL introduces a feedback-driven training framework that improves a model’s tool-use capabilities by leveraging a verifiable reward function, which balances tool invocation accuracy and task completion using only environmental feedback. Nemotron-Research-Tool-N1 (Tool-N1) [25] is a series of LLMs trained with R1-style reinforcement learning to enhance tool-calling capabilities in multi-tool reasoning scenarios. At each action step, the model produces explicit reasoning enclosed in <think> tags, followed by structured tool calls in <tool_call> tags, effectively separating internal reasoning from external tool execution. WebGen-Agent [89] introduced a novel framework for interactive website code generation and integrated a visual-language model for assessing website appearance via screenshots and a GUI-agent for testing functional correctness. To enhance smaller open-source models, the authors propose Step-GRPO with Screenshot and GUI-agent Feedback, a step-level reinforcement learning method that uses appearance and functionality scores as dense rewards. ToolExpander [90] enhanced GRPO-based RL for single-turn tool tasks, specifically targeting small-scale, resource-constrained LLMs. It introduced Dynamic Multi-Round Hard Sampling to replace difficult samples with high-quality few-shot examples during training, reducing the proportion of hard samples and improving learning stability. Additionally, the Self-Exemplifying Thinking mechanism allows the model to autonomously generate and analyze few-shot examples, receiving a small extra reward to encourage self-guided learning.
More tasks A1-type training is also applied to many other downstream tasks achieving promising results. Rec-R1 [91] (TMLR 2025) is a reinforcement learning framework that directly optimizes LLMs for recommendation tasks using feedback from downstream recommendation systems, rather than relying on imitation of other models or synthetic supervised data. By casting LLM generation as a policy and using recommendation metrics (e.g., NDCG, Recall) as reward signals, Rec-R1 enables closed-loop adaptation of LLM outputs, aligning generation with actual recommendation performance. SQL-R1 [92] is a recent work that addresses the Natural Language to SQL (NL2SQL) task by leveraging reinforcement learning to enhance reasoning capabilities in complex database scenarios. SQL-R1 introduces a tailored RL reward function comprising format, execution, result, and length rewards to guide the model toward generating SQL queries that accurately reflect user intent. olmOCR 2 [93] is a state-of-the-art open-source OCR system for converting digitized print documents into naturally ordered plain text. This reinforcement learning-based method is a key departure from the project’s first version, olmOCR 1 [94], which was trained using supervised fine-tuning to mimic the static outputs of a teacher model. Critically, this new approach moves beyond traditional edit distance metrics, which often fail to capture practical correctness or handle layout ambiguities. Instead, the model is trained using a diverse set of binary unit tests as the reward signal, covering text presence/absence, reading order, table accuracy, and math formula rendering.
In summary, RLVR-based A1 methods represent a substantial evolution, directly engaging with interactive environments to iteratively improve performance. RLVR-based A1 methods leverage environment-derived reward signals to guide policy optimization, often integrating advanced techniques such as KL-regularized PPO, GRPO, and dynamic sampling. The unifying principle is that models learn through trial-and-error, receiving immediate feedback from the environment to refine both reasoning and tool-use strategies. Despite their effectiveness, RLVR-based approaches generally require careful reward design, computational resources for interactive training, and mechanisms to stabilize learning.
@@ 4.2 A2: Agent Output as Signal
Different from A1-type adaptation, which leverages feedback obtained from tool executions or external environments, the A2 paradigm focuses on using evaluations of the agent’s own outputs as the optimization signal. In this setting, the learning or adjustment process is driven by assessing the quality of the agent’s generated outputs. Such evaluations may come from human judgments, automated metrics, or environment-based rewards, and are used to update or refine the agent policy. Under this paradigm, adaptation can occur in two primary settings:
  • Agent Adaptation w/o Tools: The agent relies on the evaluation of its reasoning or problem-solving outputs without involving external tools. This direction mainly focuses on improving intrinsic reasoning abilities, such as mathematical reasoning, coding, or logical inference, by optimizing the model based on evaluations of its generated solutions.
  • Agent Adaptation w/ Tools: The agent’s generated outputs are assessed in conjunction with tool interactions, providing feedback on how effectively the agent plans, selects, and executes tool usage. This line of adaptation aims to enhance the agent’s capability in coordinating and utilizing tools, using evaluation signals derived from task outcomes that depend on tool-mediated actions.
Table 2: A2 Methods: Tool Adaptation w/ Agent Supervision
@@ 4.2.1 Agent Adaptation w/o Tools
A major breakthrough in the paradigm of output-based agent adaptation emerged with the introduction of the DeepSeek-R1 framework [24] (Nature 2025), which demonstrated that reinforcement learning with verifiable reward (RLVR) can effectively enhance the reasoning capabilities of large agents. In this framework, a strong base model serves as the core reasoning engine, while reinforcement learning encourages the generation of outputs that are logically consistent and verifiably correct. The training process focuses primarily on reasoning-intensive domains such as mathematics and code generation, where the quality of outputs can be automatically evaluated through deterministic correctness signals. This approach not only improved model reasoning but also revealed a scalable pathway for further enhancing agent intelligence beyond supervised fine-tuning.
Concurrently, Kimi-1.5 [95] advanced this paradigm by scaling reinforcement learning for multi-modal agents and introducing simplified yet effective policy optimization strategies. By leveraging large-scale reasoning data and efficient reward modeling, Kimi-1.5 achieved strong performance across a range of reasoning benchmarks, matching or surpassing prior state-of-the-art models. Together, these works sparked a new wave of research into the so-called R1 paradigm, where reinforcement learning is used to refine the reasoning process of agentic systems based on verifiable output evaluations. Following this development, a number of subsequent studies have extended the paradigm to various applications and reasoning settings.
Following the emergence of the R1 paradigm, a series of subsequent works further extended the idea of optimizing agent reasoning through evaluations of model outputs, without relying on external tools. These studies explored diverse learning signals, objectives, and task domains, collectively enriching the landscape of output-based adaptation for reasoning enhancement.
Empower [96] proposed a self-supervised fine-tuning framework for assistive language models, where the optimization objective centers on maximizing human empowerment rather than explicit correctness. By using only offline text data, the method aligns agents to assist human users in multi-turn coding tasks, encouraging context-sensitive and cooperative behavior without the need for additional feedback or verifiable rewards.
KnowRL [97] introduced a reinforcement-based approach to strengthen self-knowledge calibration. Instead of focusing on task-specific correctness, KnowRL trains agents to assess their own confidence and feasibility in producing reliable answers. Through internally generated rewards derived from self-assessment, the model enhances its awareness of what it knows and what it does not, improving reliability and consistency across reasoning domains.
GRACE [98] reimagines contrastive learning as a form of reward-guided optimization, transforming contrastive objectives into policy signals that encourage explicit, interpretable reasoning. By treating positive–negative sample distinctions as reward feedback, GRACE bridges generative reasoning and representation learning, yielding improved embedding alignment and more transparent rationales. A related study, Rec-R1 [91], applies reinforcement optimization to product re-ranking tasks, demonstrating that reinforcement signals derived from task-specific output evaluations can improve discriminative performance while preserving general reasoning ability. EHRMind [99] extends the reinforcement-with-verifiable-reward framework to clinical reasoning scenarios. Targeting electronic health record (EHR) interpretation, the study highlights the limitations of RLVR alone, noting that domain-specific reasoning often requires prior knowledge alignment through SFT. EHRMind combines a lightweight SFT warm-up phase with subsequent RLVR optimization, effectively stabilizing training and improving interpretability in medical tasks such as clinical calculation, patient-trial matching, and disease diagnosis. This finding underscores a broader insight: while RL-based adaptation enhances reasoning quality, SFT remains a critical foundation for domain grounding and stable agent adaptation.
Before the emergence of the R1 paradigm, several studies had already explored output-based adaptation strategies that optimize reasoning quality through self-generated feedback, without modifying model parameters via traditional supervised signals. These early attempts laid the conceptual foundation for reinforcement-driven reasoning improvement later seen in R1-style frameworks.
Self-Refine [100] (NeurIPS 2023) introduced an iterative refinement framework in which the same language model acts as both generator and critic. The model first produces an initial response and then evaluates and revises it based on self-generated textual feedback. This process, inspired by human-style revision, improves output quality across diverse domains such as dialogue, mathematics, and code generation. Notably, Self-Refine requires no supervised data, auxiliary models, or reinforcement learning, which shows that structured self-feedback alone can lead to measurable gains in reasoning accuracy and output preference.
Building upon this direction, SCoRe [100] (ICLR 2025) proposed a reinforcement learning approach for enabling language models to self-correct using entirely self-generated data. Unlike conventional SFT, which struggles to teach effective correction behavior, SCoRe employs multi-turn online reinforcement learning to encourage models to iteratively refine their reasoning under their own distribution of responses. By combining reward regularization and self-generated correction traces, the method significantly improves self-correction ability on mathematical and reasoning benchmarks, demonstrating that reinforcement learning can effectively operationalize self-reflection into a stable learning signal.
Backpropagating language-model feedback. TextGrad [101] (Nature 2025) introduced a general framework for agent self-improvement through textual gradient descent (TGD). Instead of relying on numerical gradients or verifiable environment rewards, TextGrad propagates language-model feedback in the form of natural-language critiques that describe how to improve the model’s outputs. These feedback messages act as “textual gradients,” allowing optimization across black-box LLM systems without requiring access to their internal parameters. This method formalizes self-refinement as a differentiable-like process and generalizes earlier outcome-based approaches such as DeepSeek-R1. Empirically, TextGrad improves GPT-4o’s zero-shot code accuracy on LeetCode-Hard (from 26% to 36%), raises MMLU-Physics performance from 91.2% to 95.1%, and enhances the multi-tool agent Chameleon by 7.7%. Conceptually, TextGrad exemplifies the A2 paradigm, extending reinforcement-style adaptation from scalar rewards to structured linguistic feedback. It unifies prompt tuning, reasoning refinement, and compound-agent optimization under a single abstraction of backpropagating language feedback, marking a step toward interpretable and parameter-agnostic agent adaptation. Building on this, metaTextGrad [102] (NeurIPS 2025) applies the A2 paradigm recursively to the optimizer itself, using validation feedback to automatically refine the optimizer’s prompts and structure for better task alignment.
@@ 4.2.2 Agent Adaptation w/ Tools
Following the rise of the R1 paradigm, the idea of using agent outputs as optimization signals has expanded beyond pure reasoning tasks to encompass tool-using agents. These studies extend output-based RL to settings where agents must decide when and how to invoke external tools, integrating real-time feedback from retrieval systems, APIs, or executable environments. This marks a shift from reasoning-centric refinement to tool-grounded adaptation, where outcome feedback from the external world provides rich and verifiable learning signals.
Retrieval-based tool learning. Different from earlier work like Self-RAG [103] (ICLR 2024) and its successors [104, 105, 106], which introduces an distillation-SFT-based paradigm to teach models to use retrieval tools for search, a recent major line of work investigates how RL can improve the use of retrieval tools for question answering. R1-Searcher [107], Search-R1 [49] (COLM 2025), ReSearch [108] (NeurIPS 2025), and their successors [109, 110, 111, 112, 113, 114, 115] all extend the R1 paradigm by enabling LLMs to autonomously generate and refine search queries during multi-turn reasoning. R1-Searcher [107] proposes a two-stage RL framework that incentivizes the use of external search APIs, enhancing factual accuracy and reducing hallucinations in open-domain QA. Trained with multi-turn RL, the model learns to balance reasoning and retrieval, achieving up to 24% improvement over strong RAG baselines. Similarly, Search-R1 [49] formulates search invocation as a reinforcement optimization problem, where retrieved evidence and final correctness jointly form the outcome-based reward. ReSearch [108] trains LLMs to reason with search via reinforcement learning, without any supervised data on reasoning steps. It integrates search queries and retrieved results directly into the reasoning chain using tags such as <think>, <search>, and <result> and optimizes the model with GRPO to decide when and how to search. Trained on multi-hop QA tasks, ReSearch yields 9–22% absolute gains over iterative RAG baselines and exhibits emergent reflection and self-correction behaviors during RL training These works demonstrate that reasoning-oriented RL can naturally extend to retrieval-augmented contexts, allowing LLMs to internalize when external information is needed.
Code- and execution-based tool learning. A parallel direction focuses on code-based environments where tools provide executable feedback. CodePRM [116] (ACL 2025) introduces a process reward model that scores reasoning steps based on code execution results, forming a Generate–Verify–Refine pipeline that dynamically corrects reasoning errors during inference. ReTool [50] advances this idea by integrating real-time code execution into RL rollouts, teaching models when and how to invoke computational tools such as interpreters to optimize mathematical and symbolic reasoning.
@@
General multi-tool and agentic learning. Beyond retrieval, several studies generalize this principle to agents interacting with diverse APIs and environments. Test-Time Self-Improvement (TT-SI) [117] introduces on-the-fly self-improvement, where the agent identifies uncertain test cases and generates new training data for them, performing fine-tuning directly at inference time. Agent Lightning [118] provides a flexible RL framework that decouples agent execution from training, allowing reinforcement optimization to handle complex, multi-agent, and multi-tool workflows with minimal code modification. Re-ReST [119] extends self-training with reflection, using environment feedback such as unit test results to refine low-quality trajectories, yielding large gains on HotpotQA and AlfWorld. In a similar spirit, Self-Challenging Agents [120] introduce a self-generated curriculum: the model first generates novel tool-use tasks as a challenger, then solves them via RL as an executor, achieving more than a twofold improvement on multi-turn tool-use benchmarks. Agent-R [121] further formalizes iterative self-reflection via model-guided critique construction, continuously correcting failed trajectories using Monte Carlo Tree Search (MCTS) rollouts, which improves performance by 5.6% across interactive environments. Finally, A2FM [122] unifies reasoning and acting within a cost-regularized RL framework, dynamically choosing between internal reasoning, tool invocation, or direct answering, thereby improving both efficiency and accuracy in hybrid reasoning benchmarks. Addressing system-level scalability in a similar vein, VerlTool [123] introduces a unified infrastructure for agentic RL. By decoupling the RL workflow from a modular tool server and implementing asynchronous rollouts, it eliminates synchronization bottlenecks, enabling the efficient training of multi-modal agents across disparate tasks ranging from visual reasoning to software engineering.

### v3

@@ 4.1.2 RLVR-Based Methods
Formal theorem proving [69, 70, 71, 72, 73, 74, 75, 76, 77] provides a canonical domain for RLVR under the A1 paradigm, as proof assistants provide ground-truth, tool-execution-signaled feedback at every step. In this setting, the agent proposes one or more tactics (i.e., proof steps), a formal proof checker (the tool) deterministically verifies their validity, and the resulting validated proof-state transition is returned to the agent. The verification outcome (e.g., whether a tactic is accepted, whether it advances the proof state, or whether a complete proof is achieved) serves directly as a verifiable reward signal for policy optimization. Compared to code-execution RLVR, where unit tests may be sparse or incomplete, theorem proving offers step-wise semantic verification with minimal ambiguity, enabling denser rewards and substantially easing long-horizon credit assignment. Recent systems such as AlphaProof [78] (Nature 2025), DeepSeek-Prover-V2 [79] (ICLR 2025), Kimina-Prover [73], and Leanabell-Prover-V2 [80] use this verifier feedback to train multi-step proof search policies via reinforcement learning, while a complementary line of work augments the native proof checker feedback with auxiliary guidance signals to prioritize trajectories, shape exploration, or stabilize optimization on top of verifier-grounded rewards [81, 82, 83, 84, 85]. While RLVR is well suited for learning proof strategies under a fixed prover snapshot, formal theorem proving also highlights a broader adaptation challenge: formal libraries (e.g., mathlib [86]) and large, actively evolving formalization projects [87] built by the Lean community grow continuously, expanding the available premise space. Addressing this non-stationarity often requires complementary continual or low-resource adaptation mechanisms beyond pure RLVR, which we discuss in §9.2.
Multi-tool reasoning systems. When multiple tools must be composed sequentially or conditionally, the agent faces a combinatorial action space and sparser episode-level rewards. Methods in this category address the challenge through three complementary strategies: (i) routing, (ii) environment construction, and (iii) dense step-level rewards.
Router-R1 [88] (NeurIPS 2025) exemplifies routing: a policy LLM learns to alternate between internal reasoning and external model selection, dynamically invoking different LLMs from a routing pool. FTRL [89] tackles environment construction by automatically synthesizing diverse tool-use training environments through a multi-stage pipeline, so the agent can train without hand-crafted external toolsets; a verifiable reward balances invocation accuracy with task completion. Tool-N1 [26] separates internal reasoning (<think> tags) from external execution (<tool_call> tags), giving the RL optimizer distinct credit-assignment channels for thought and action. WebGen-Agent [90] demonstrates the dense-reward strategy: appearance scores from a visual-language model and functionality scores from a GUI-agent are combined via Step-GRPO to supervise interactive website code generation at every turn. ToolExpander [91] targets the resource-constrained regime with Dynamic Multi-Round Hard Sampling (replacing hard samples with few-shot exemplars) and a Self-Exemplifying Thinking mechanism that rewards self-generated analysis, stabilizing GRPO for small LLMs.
Domain-specific extensions. A1-type RLVR has also been applied beyond the core domains above. Rec-R1 [92] (TMLR 2025) casts recommendation as an RL policy over LLM generation, using NDCG and Recall as rewards. SQL-R1 [93] uses a composite reward (format, execution, result, and length) for NL-to-SQL. olmOCR 2 [94] replaces the SFT-based teacher imitation of olmOCR 1 [95] with diverse binary unit tests (text presence, reading order, table accuracy, formula rendering) as the reward signal for document OCR. These applications reinforce a common pattern: whenever a domain provides a cheap, deterministic correctness oracle, A1-style RLVR can yield large gains with modest engineering effort.
Cross-domain principles.
Despite the diversity of domains (web search, code, theorem proving, multi-tool reasoning, and specialized applications), four recurring patterns cut across all A1 RLVR work:
  • Signal density determines learning efficiency. Domains with dense, per-step feedback (theorem proving, code execution) enable faster convergence than those with sparse, episode-level rewards (multi-tool reasoning). Dense feedback explains why A1 methods in code and proof domains often require less data than A2 counterparts.
  • Reward quality dominates reward quantity. Code-R1 [66] and DeepRetrieval [22] both show that investing in clean, verified reward signals yields larger gains than scaling training data, a finding consistent across retrieval, code, and SQL domains.
  • Format rewards are necessary but not sufficient. Nearly all successful RLVR methods combine a task-specific reward with a format compliance reward (e.g., DeepRetrieval’s $r_{\text{format}}$, Tool-N1’s structured output tags). Format rewards alone cannot drive meaningful adaptation, but their absence leads to degenerate outputs.
  • Stabilization mechanisms are domain-agnostic. KL regularization, dynamic sampling, and curriculum scheduling appear across all domains, suggesting that the challenge of stabilizing on-policy RL for language agents is fundamental rather than domain-specific.
These principles notwithstanding, RLVR remains expensive: it requires careful reward design, interactive compute, and explicit stabilization—costs that motivate the tool-centric adaptation paradigms discussed in later sections.
@@ 4.2 A2: Agent Output as Signal
Whereas A1 optimizes the mechanics of individual tool calls, A2 optimizes the agent’s end-to-end output—reasoning chains, final answers, or generated artifacts—using evaluations of those outputs as the training signal. The shift from action-level to output-level feedback brings two consequences: (i) the agent can learn strategic behaviors (when to invoke a tool, how deeply to search, whether to self-correct) that A1 signals cannot reward, and (ii) credit assignment becomes harder because a single episode-level reward must be attributed across many internal decisions.
We distinguish two settings based on whether tools participate in the evaluation loop:
  • Agent Adaptation w/o Tools: The agent improves intrinsic reasoning (mathematics, coding, logical inference) by optimizing against evaluations of its generated solutions. The design space here is defined by the granularity of the evaluation signal (scalar correctness vs. structured linguistic critique) and whether weights are updated (RL-based vs. inference-time refinement).
  • Agent Adaptation w/ Tools: The agent’s outputs are assessed in conjunction with tool interactions, so the training signal reflects both reasoning quality and tool-coordination skill. The key additional challenge is learning the meta-policy—when and how to invoke tools—alongside the reasoning policy.
Table 2: A2 Methods: Agent Output Signaled Agent Adaptation
@@ 4.2.1 Agent Adaptation w/o Tools
Three design axes organize the methods below: (1) the reward granularity ranges from binary final-answer correctness to structured natural-language critiques; (2) the optimization locus ranges from weight updates (RL, SFT) to inference-time refinement (no weight changes); and (3) the objective scope ranges from task-specific correctness to broader goals such as calibration or human empowerment.
Scalar-reward RL (the R1 paradigm). DeepSeek-R1 [25] (Nature 2025) demonstrated that RLVR with binary final-answer correctness is sufficient to unlock strong mathematical and coding reasoning in large agents, revealing a scalable pathway beyond SFT. Kimi-1.5 [96] extended this to multi-modal agents with simplified policy optimization, matching or surpassing DeepSeek-R1 across reasoning benchmarks. Together, these works established the R1 paradigm: RL on verifiable output evaluations as the primary driver of reasoning improvement.
Subsequent R1-style studies explore the boundaries of this paradigm by varying the objective: EHRMind [97] applies RLVR to clinical reasoning (EHR interpretation), finding that a lightweight SFT warm-up is necessary before RL to ground domain knowledge—a pattern that recurs across specialized domains. KnowRL [98] replaces task correctness with self-knowledge calibration: the agent is rewarded for accurately assessing its own confidence, improving reliability without targeting any single task. Empower [99] goes further by defining the objective as maximizing human empowerment rather than correctness, aligning assistive agents to multi-turn coding tasks using only offline text data. GRACE [100] transforms contrastive-learning objectives into RL-style policy signals, bridging generative reasoning with representation learning. A related study, Rec-R1 [92], applies RL to product re-ranking using recommendation metrics (NDCG, Recall) as rewards. Rec-R1 straddles the A1/A2 boundary—A1 when the reward is computed from the tool’s ranking output, A2 when it evaluates the agent’s overall recommendation quality—illustrating that the distinction is not always binary but depends on where in the agent-tool pipeline the evaluation signal is computed.
Inference-time self-refinement (no weight updates). A complementary line of work achieves adaptation at inference time, avoiding the cost of retraining. Self-Refine [101] (NeurIPS 2023) introduced the generate-critique-revise loop: the same LLM produces an initial response, evaluates it, and revises it based on self-generated textual feedback, improving output quality across dialogue, mathematics, and code without any supervised data or auxiliary models. SCoRe [101] (ICLR 2025) makes self-correction trainable: multi-turn on-policy RL encourages the model to iteratively refine its own responses, operationalizing the Self-Refine loop into a stable learning signal. The progression from Self-Refine to SCoRe mirrors the SFT-to-RLVR progression in A1: inference-time heuristics are replaced by learned policies that improve with scale.
Structured linguistic feedback. TextGrad [102] (Nature 2025) generalizes the scalar reward to natural-language critiques that describe how to improve the output. These “textual gradients” enable optimization across black-box LLM systems without parameter access, improving GPT-4o’s zero-shot code accuracy on LeetCode-Hard from 26% to 36% and MMLU-Physics from 91.2% to 95.1%. metaTextGrad [103] (NeurIPS 2025) applies the same principle recursively to the optimizer itself, using validation feedback to refine the optimizer’s prompts. TextGrad and metaTextGrad occupy a unique position: unlike RL methods, they require no reward function design; unlike Self-Refine, they propagate structured credit through multi-component systems.
Synthesis of A2 without tools.
The methods above map onto a two-dimensional design space. Along the first axis (reward granularity), scalar final-answer correctness (DeepSeek-R1) is broadly applicable but sparse, while structured linguistic feedback (TextGrad) is information-rich but requires a capable critic model. Along the second axis (optimization locus), weight-update methods (RL, SFT) amortize adaptation cost across future queries, while inference-time methods (Self-Refine, TextGrad) pay per-query cost but require no training infrastructure. A recurring finding across both axes is that SFT warm-up followed by RL yields more stable training than RL alone: SFT provides domain grounding, while RL refines the policy toward the evaluation objective.
@@ 4.2.2 Agent Adaptation w/ Tools
When tools enter the A2 loop, the agent must learn a meta-policy: not only how to reason but when to invoke a tool, which tool to choose, and how to integrate tool outputs back into the reasoning chain. The reward is still computed on the agent’s final output, but the policy space is much larger, making credit assignment and data efficiency the central challenges.
Retrieval-based tool learning. Prior distillation-SFT approaches such as Self-RAG [104] (ICLR 2024) and its successors [105, 106, 107] teach retrieval tool use from expert demonstrations; RL-based methods instead let the agent discover search strategies on its own. R1-Searcher [108], Search-R1 [50] (COLM 2025), ReSearch [109] (NeurIPS 2025), and their successors [110, 111, 112, 113, 114, 115, 116] all train LLMs to autonomously generate and refine search queries during multi-turn reasoning. The distinguishing design choice across this family is the action representation: R1-Searcher incentivizes search API calls via a two-stage RL framework (up to 24% over RAG baselines), Search-R1 formulates search invocation as a joint reward over retrieved evidence and final correctness, and ReSearch integrates queries and results directly into the reasoning chain via <think>/<search>/<result> tags, optimized with GRPO. ReSearch is notable for exhibiting emergent reflection and self-correction behaviors—strategic capabilities that were not explicitly supervised but arose from holistic task optimization.
Code- and execution-based tool learning. Code execution provides a natural bridge between A1 and A2: the execution outcome is verifiable (A1-like), but the training signal evaluates the agent’s overall problem-solving trajectory (A2-like). CodePRM [117] (ACL 2025) scores intermediate reasoning steps via a process reward model trained on code execution results, forming a Generate-Verify-Refine pipeline that corrects errors during inference. ReTool [51] integrates real-time code execution directly into RL rollouts, so the agent learns when to invoke an interpreter to offload computation—illustrating how A2 training can discover tool-use strategies that A1 training cannot, because the reward depends on whether the agent’s overall answer (not just the tool call) is correct.
General multi-tool and agentic learning. In the most general setting, agents interact with heterogeneous APIs and environments. Three recurring themes organize the methods below: data generation, self-reflection, and infrastructure.
@@
On the data side, Self-Challenging Agents [118] introduce a self-generated curriculum—the model first creates tool-use tasks as a challenger, then solves them via RL as an executor—achieving over twofold improvement on multi-turn benchmarks. Re-ReST [119] uses environment feedback (e.g., unit-test results) to refine low-quality trajectories through reflection, yielding large gains on HotpotQA and AlfWorld. On the self-reflection side, Agent-R [120] formalizes iterative self-correction via model-guided critique and MCTS rollouts (+5.6% across interactive environments), and Test-Time Self-Improvement [121] performs on-the-fly fine-tuning at inference time on self-identified uncertain cases. On the infrastructure side, Agent Lightning [122] decouples agent execution from training to support complex multi-agent workflows, A2FM [123] unifies reasoning and acting under cost-regularized RL (dynamically choosing between thinking, tool use, and direct answering), and VerlTool [124] provides asynchronous rollout infrastructure that eliminates synchronization bottlenecks in agentic RL.
Synthesis of A2 with tools.
A2 methods with tools share a defining characteristic: the agent must learn not only how to use tools (also targeted by A1) but when to use them and how to fold their outputs into a coherent reasoning chain. This strategic dimension explains why A2 training frequently produces emergent behaviors—self-correction, adaptive search depth, query refinement—that are never explicitly supervised. The cost is data efficiency: the sparse, episode-level reward requires either large training sets (e.g., Search-R1’s $\sim$170k examples) or carefully designed curricula (e.g., Self-Challenging Agents’ self-generated tasks). The tension between A1’s per-action precision and A2’s end-to-end strategic optimization is analyzed quantitatively in §6.